A 50 kg child slides down a 35° with a µ of -0.2 and length of 10 meters. What is the kinetic energy the child has at the bottom of the slide due to friction? What was the child’s gravitational potential energy at the top of the slide (Remember the slide is on an angle, the child’s height is not 10 meters)?


1
Expert's answer
2021-05-17T17:34:36-0400
"U=mgl\sin{35}\\U=(50)(9.8)(10)\sin{35}=2810.5\ J"

"K=mgl\sin{35}-\mu mgl\cos{35}\\K=(50)(9.8)(10)(\sin{35}-0.2\cos{35})=2007.8\ J"